Simon is rolling a fair number cube and flipping a coin.What is the probability that he will roll the number 3 and the coin will land on heads?
Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

Simon is rolling a fair number cube and flipping a coin.What is the probability that he will roll the number 3 and the coin will land on heads?
--------
The events are independent so you should multiply the separate probabilities.
P(rolling a 3)=1/6
P(heads)=1/2
P(roll 3 and heads)=1/6*1/2=1/12
Cheers,
stan H.

Answer by checkley71(8403)

THE DIE HAS A 1/6 CHANCE OF ROLLING A 3 & THE COIN HAS A 1/2 CHANCE OF BEING A HEAD. THUS THE COMBINED PROBABILITY OF A 3 & A HEAD IS 1/6*1/2=1/12.
IOF YOU WANT TO PROVE THIS SET UP A TABLE WITH ALL POSSIBLE OUTCOMES
H & 1,2,3,4,5,6 & T & 1,2,3,4,5,6. YOU'LL FIND THERE ARE 12 OUTCOMES & ONLY ONE IS A 3 & A HEAD.